### 10 Mar 2026 v1

@@ 1 Introduction
In this paper we focus on two harvesting mechanisms in age-structured models. The first is a rate-control formulation, in which harvesting enters the state equation as a direct removal term. The second is an effort-control formulation, in which harvesting acts as an additional mortality intensity and the instantaneous yield is proportional to both effort and stock abundance. First, for the infinite-horizon harvesting problem, we derive necessary optimality conditions of Pontryagin type, including the adjoint system, transversality condition, switching relations for the distributed harvesting control and the boundary inflow control, and complementary slackness for the state constraint. We then compare rate-control and effort-control formulations to highlight the structural differences between them. In particular, we exhibit the nonlocal coupling term in the effort-control adjoint equation induced by aggregate population dependence and thereby make explicit how the choice of harvesting mechanism changes the analytical structure of the optimality system.
@@
The contributions of the paper can be summarized as follows:
@@
Figure 1 provides a mechanistic map of the paper. The paper is organized as follows. Section 2 reviews the relevant literature on age-structured population dynamics, steady-state analysis, optimal control, and bioeconomic harvesting models. In Section 3 we introduce the models and standing assumptions. In Section 4 we formulate the infinite-horizon optimal control problem and derive necessary optimality conditions. Section 5 discusses the effort-control formulation and the associated formal adjoint equation with nonlocal dependence. Section  6 collects autonomous reductions and explicit stationary formulas, and Section 7 concludes with a discussion of the mathematical implications of the two harvesting mechanisms and possible directions for future work. Appendix A presents numerical simulations.
@@ 2 Literature review
The present paper lies at the intersection of five closely related strands of research: age-structured population dynamics, steady-state analysis, optimal control of structured populations, infinite-horizon maximum principles, and bioeconomic harvesting models.
2.1 Age-structured population dynamics
The continuous-time theory of age-structured populations begins with the McKendrick–von Foerster equation, which has become a standard framework for demographic and ecological dynamics [mkendrick_applications_1925, von_foerster_remarks_1959]. A systematic mathematical treatment of nonlinear age-dependent population models is developed by Webb [webb_semigroup_1984, webb_theory_1985], while Iannelli [iannelli_mathematical_1995] and Iannelli and Milner [iannelli_basic_2017] provide rigorous introductions to the analytic and numerical theory, respectively. The broader structured-population perspective, including age and physiological structure, is represented by Metz et al. [metz_dynamics_1986] and Cushing [cushing_introduction_1998]. The semigroup and abstract-evolution approaches may be found in [engel_one-parameter_2000, banasiak_perturbations_2006, magal_theory_2018]. More recent developments on structured models with nonlocal feedback, stability, and lineage-type organization may be found in [wang_damage-structured_2026, liang_global_2025], where feedback-regulated structured dynamics are analyzed in biologically distinct settings.
2.2 Steady states analysis
The analysis of steady states in age-structured models goes back to the nonlinear framework of Gurtin and MacCamy [gurtin_non-linear_1974, gurtin_simple_1979] and the implicit function methods of Cushing [cushing_existence_1984]. Scalar and operator-theoretic reductions of steady-state problems were developed by Prüss [pruss_equilibrium_1981, pruss_stability_1983] and by Busenberg et al. [busenberg_endemic_1988]. A general formulation of nonlinear structured steady states in terms of consistency relations and next-generation operators was given by Diekmann et al. [diekmann_steady-state_2003]. The corresponding threshold phenomena, organized around reproduction numbers, persistence, and extinction, is developed in the work of Thieme [thieme_mathematics_2018] and Inaba [inaba_age-structured_2006]; see also [walker_positive_2009, wang_algebraicspectral_2026, yu_dedifferentiation_2026] for related analyses of threshold conditions and stability.
2.3 Optimal control of age-structured populations
Optimal control for age-structured systems has been studied extensively in both finite- and infinite-horizon settings. Early Pontryagin-type formulations for age-dependent controls were developed by Brokate [brokate_pontryagins_1985], while Barbu and Iannelli [barbu_optimal_1999] established variational formulations and existence results. A comprehensive treatment is given in the monograph of Aniţa [anita_analysis_2000]; see also [anita_optimal_1998]. On the harvesting side, structural properties of optimal policies have been investigated in [hritonenko_structure_2007], while more general control-theoretic expositions may be found in [veliov_optimal_2008]. Furthermore, for stochastic ecological systems, recent analytical frameworks for well-posedness and asymptotic bounds have been developed in [wang_analysis_2025], suggesting a possible route toward stochastic extensions of structured harvesting models.
2.4 Infinite-horizon optimal control and transversality
Infinite-horizon optimal control problems require additional care because the adjoint system must be complemented by a transversality condition at infinity. Classical references include Arrow and Kurz [arrow_optimal_1969] and Carlson et al. [carlson_infinite_1991]. For general infinite-horizon maximum principles, a central role is played by the work of Aseev and Kryazhimskii [aseev_pontryagin_2004], the survey of Aseev et al. [aseev_infinite-horizon_2012], and the refinements of Aseev and Veliov [aseev_maximum_2015]. In the age-structured setting, the interaction between transport dynamics, discounting, and the adjoint system creates additional analytical difficulties; see in particular the discussion of infinite-horizon age-structured optimal control in [aseev_infinite-horizon_2012].
2.5 Rate-control, effort-control, and bioeconomic structure
From the bioeconomic point of view, the distinction between direct harvesting rates and effort-based mortality adjustment is classical. Clark [clark_economics_1973] established the basic economic framework for renewable-resource management, while Getz and Haight [getz_population_1989] and Reed [reed_optimal_1979] analyzed related harvesting mechanisms in demographic and stock-recruitment models. In age-structured settings, the difference is mathematically significant: rate-control enters additively and preserves an affine transport structure, whereas effort-control enters multiplicatively and leads to a bilinear coupling between control and state. When mortality also depends on the aggregate population size, the adjoint equation acquires a nonlocal term coupling all ages through the total stock. Related density-dependent and structured formulations appear in [boucekkine_optimal_2013].
2.6 Position of the present paper
The present paper uses the McKendrick–von Foerster framework as its basic setting, adopts an infinite-horizon variational argument, and places rate-control and effort-control formulations in direct comparison. Its main purpose is not to survey these literature exhaustively, but rather to connect specific issues within a single age-structured framework: the necessary optimality system for the infinite-horizon harvesting problem and the formal emergence of a nonlocal adjoint term in the effort-control case.
@@ 6 Autonomous reductions and explicit stationary formulas
7 Discussion
This paper presents two different formulations: rate-control and effort-control harvesting formulations. For the rate-control formulation, the associated infinite-horizon optimization problem admits a Pontryagin-type necessary optimality system. In current-value form, this system consists of the adjoint transport equation, the terminal-age and transversality conditions. Besides, we study the switching relations for the distributed control and the boundary inflow, and the complementary slackness condition associated with the state constraint $x\geq 0$.
A central contribution of the paper is the clarification of the structural divide between rate-control and effort-control mechanisms. In the rate-control model, harvesting enters additively, which leads to an affine Volterra-type stationary profile and a local adjoint equation. By contrast, effort-control acts multiplicatively. When natural mortality depends on the aggregate population size $E(t)$, the effort-control adjoint acquires the nonlocal coupling term
$\int_{0}^{A}\mu_{E}(E(t),s)x(t,s)\lambda(t,s)\,ds,$
which couples all ages through the total population level. Thus, direct removal and mortality-intensity adjustment are not merely alternative parameterizations of the same model, but generate genuinely different analytical structures.
From a methodological point of view, the paper provides a rigorous variational derivation for the rate-control optimality system, while the effort-control problem is treated at the level of a formal adjoint calculation. This distinction is deliberate: the nonlocal and nonlinear structure of the effort-control dynamics requires a separate well-posedness and differentiability theory.
Natural directions for future work include the dynamical stability of the supercritical positive equilibrium, a fully rigorous functional-analytic treatment of the infinite-horizon transversality condition, and the well-posedness of the nonlinear nonlocal effort-control state equation. First, it would be interesting to connect the present structured-population framework with multiscale hybrid models that combine continuum dynamics with agent-based components [liu_bidirectional_2025, wang_analysis_2025-1]. The hybrid framework will be a natural setting to incorporate more complex environmental factors. Second, it would also be interesting to connect the present age-structured framework with hybrid PDE-based models that combine transport, feedback, and mode-selection mechanisms in multiscale environments, as in recent studies of tumor–stroma patterning and hybrid PDE–ODE mass-transport systems [yu_chemotactic_2026, yu_mode-wise_2026]. A further direction is the incorporation of demographic fluctuations into structured harvesting models. Recent stochastic ecological work has emphasized that covariance-consistent noise and absorbing-boundary formulations can be essential for correctly capturing extinction structure [yu_full-covariance_2026, yu_beyond_2026].

### 02 Apr 2026 v2

@@ 1 Introduction
In this paper we focus on two harvesting mechanisms in age-structured models. The first is a rate-control formulation, in which harvesting enters the state equation as a direct removal term. The second is an effort-control formulation, in which harvesting acts as an additional mortality intensity and the instantaneous yield is proportional to both effort and stock abundance. First, for the infinite-horizon harvesting problem, we derive necessary optimality conditions of Pontryagin type, including the adjoint system, transversality condition, switching relations for the distributed harvesting control and the boundary inflow control, and complementary slackness for the state constraint. We then compare rate-control and effort-control formulations to highlight the structural differences between them. In particular, we exhibit the nonlocal coupling term in the effort-control adjoint equation induced by aggregate population dependence and thereby make explicit how the choice of harvesting mechanism changes the analytical structure of the optimality system. The contributions of the paper can be summarized as follows:
@@
Figure 1 provides a mechanistic map of the paper. The paper is organized as follows. Section 2 reviews the relevant literature on age-structured population dynamics, steady-state analysis, optimal control, and bioeconomic harvesting models. In Section 3 we introduce the models and standing assumptions. In Section 4 we formulate the infinite-horizon optimal control problem and derive the necessary optimality conditions. Section 5 discusses the effort-control formulation and the associated formal adjoint equation with nonlocal dependence. Section 6 collects autonomous reductions and explicit stationary formulas, and Section A presents numerical simulations. Section 7 concludes with a discussion of the mathematical implications of the two harvesting mechanisms and possible directions for future work.
@@ 2 Literature review
The present paper lies at the intersection of age-structured population dynamics, steady-state analysis, optimal control, infinite-horizon maximum principles, and bioeconomic harvesting models.
The continuous-time theory of age-structured populations is rooted in the McKendrick–von Foerster equation [mkendrick_applications_1925, von_foerster_remarks_1959]. This foundational framework has been systematically expanded through nonlinear, semigroup, and abstract-evolution approaches [webb_semigroup_1984, webb_theory_1985, iannelli_mathematical_1995, iannelli_basic_2017, metz_dynamics_1986, cushing_introduction_1998, engel_one-parameter_2000, banasiak_perturbations_2006, magal_theory_2018], including recent advances in models with nonlocal feedback and physiological structure [wang_damage-structured_2026, liang_global_2025]. Parallel to demographic modeling, the analysis of steady states and related topics has evolved from early nonlinear frameworks and implicit function methods [gurtin_non-linear_1974, gurtin_simple_1979, cushing_existence_1984, pruss_equilibrium_1981, pruss_stability_1983, busenberg_endemic_1988, diekmann_steady-state_2003] to general formulations of threshold conditions, reproduction numbers, and stability [thieme_mathematics_2018, liu_bidirectional_2025, inaba_age-structured_2006, wang_analysis_2025-1, walker_positive_2009, wang_algebraicspectral_2026, yu_pattern_2026].
In the realm of resource management, optimal control for age-structured systems has been extensively developed across both finite- and infinite-horizon settings [brokate_pontryagins_1985, barbu_optimal_1999, anita_analysis_2000, anita_optimal_1998]. Significant attentions are given to the structural properties of optimal policies [hritonenko_structure_2007, veliov_optimal_2008] and recent pathways toward stochastic extensions [wang_analysis_2025]. Infinite-horizon problems introduce distinct mathematical challenges—particularly the necessity of transversality conditions to complement the adjoint system—which have been rigorously addressed in optimal control literature [arrow_optimal_1969, carlson_infinite_1991, aseev_pontryagin_2004, aseev_infinite-horizon_2012, aseev_maximum_2015]. Bioeconomically, the structural distinction between direct harvesting rates and effort-based mortality is classical [clark_economics_1973, getz_population_1989, reed_optimal_1979]. Mathematically, rate-control preserves an affine transport structure, whereas effort-control introduces a bilinear coupling, often complicated by density-dependent nonlocal terms [boucekkine_optimal_2013]. Building on these foundations, the present paper uses the McKendrick–von Foerster framework to directly compare these harvesting formulations, specifically connecting the infinite-horizon necessary optimality system with the formal emergence of a nonlocal adjoint term in the effort-control case.
@@ 6 Autonomous reductions and explicit stationary formulas
7 Discussion and Conclusion
In this paper, we have systematically compared two fundamental harvesting mechanisms within the framework of continuous-time, age-structured population dynamics. By juxtaposing a rate-control formulation with an effort-control formulation, we have highlighted the profound ways in which the choice of management variable dictates the analytical and structural properties of the resulting optimal control problem.
Mathematical and Structural Implications
From a mathematical perspective, the distinction between additive and multiplicative controls is stark. In the rate-control setting, the control variable $u(t,a)$ enters the state equation as a direct, additive sink. This preserves an affine transport structure, allowing for the rigorous derivation of an infinite-horizon maximum principle. The necessary optimality conditions, derived in Section 4, reveal an adjoint equation that is strictly local in the age variable, driven only by the discount rate, natural mortality, and the state-constraint multiplier.
Conversely, the effort-control formulation introduces a bilinear coupling through the multiplicative term $w(t,a)x(t,a)$. More importantly, incorporating a density-dependent mortality coefficient $\mu(E(t),a)$ transforms the system into a nonlinear transport equation. As demonstrated in Section 5, this aggregate population dependence fundamentally alters the dual system. The variational derivation exposes a nonlocal integral term in the adjoint equation,
$\int_{0}^{A}\mu_{E}(E(t),s)x(t,s)\lambda(t,s)\,ds,$ (58)
which formally couples the shadow prices across all ages through the total stock. This nonlocal feedback presents significant analytical challenges but offers a mathematically richer representation of density-dependent ecological constraints.
Bioeconomic and Ecological Insights
The analytical differences between the two models are closely mirrored in their bioeconomic behavior, as evidenced by our stationary reductions and numerical simulations. The rate-control mechanism is mathematically aggressive; the explicit affine Volterra representation shows that direct removal can rapidly drive specific age classes to zero. This requires strict nonnegativity constraints and heavily relies on bang-bang control strategies to prevent localized population collapse. The corresponding yield curve grows steeply but abruptly levels off as state constraints become active.
By contrast, the effort-control mechanism is inherently self-regulating. The explicit multiplicative exponential profile ensures that the extraction scales proportionally with the available stock. Consequently, the yield curve is strictly concave, and population depletion is less severe at equivalent nominal harvesting intensities. This mathematical behavior aligns with classical bioeconomic intuition: effort-based management naturally cushions against overexploitation because diminishing stock abundance simultaneously reduces the instantaneous catch.
Limitations and Future Directions
While our results clarify the structural dichotomy between rate and effort controls, several avenues for future research remain. First, the derivation of the adjoint system for Model E presented here is primarily formal. Establishing a rigorous, functional-analytic proof of the infinite-horizon maximum principle for this nonlinear, nonlocally coupled system—particularly concerning the precise formulation of the transversality conditions at infinity—remains a formidable mathematical challenge.
Second, our framework assumes purely deterministic population dynamics. Integrating stochastic demographic fluctuations or environmental noise into these structured models is a critical next step. Recent advances in the well-posedness of stochastic ecological systems suggest a viable pathway for generalizing the present optimal harvesting results to noisy, uncertain environments.
Finally, extending the current single-species framework to multi-species models with age-structured predator-prey dynamics or spatial heterogeneity would allow for the investigation of optimal harvesting policies under complex ecological interactions. The nonlocal adjoint structures identified in this paper will likely play a central role in resolving the optimality systems of such high-dimensional bioeconomic models.